# Dynamical Formation of Apparent Horizons due to Boundary Effect in Vacuum Einstein Gravity

### v2

@@ 3. Gluing of Cauchy data and horizon exclusion
(a) Semi–Global Existence. For every sufficiently large fixed constant $\mathcal{I}>0$, there exists $a_{0}=a_{0}(\mathcal{I})\gg 1$ such that the following holds. Let $a>a_{0}$ and let $N\in\mathbb{N}$ be sufficiently large. Suppose the initial data set is prescribed along the characteristic hypersurfaces $H_{u_{\infty}}$ and $\underline{H}_{0}$, where $u_{\infty}<0$ is large in magnitude, and satisfies the boundedness property of the norms described in 37-42
(43) $\Gamma^{0}_{2}+\Gamma^{0}_{\underline{\hat{\chi}}2}+\mathcal{R}^{0}_{2}+\sup_{\underline{u}}\mathcal{R}(u_{\infty},\underline{u})+\sup_{u}\mathcal{R}(u,0)<\mathcal{I}^{0},~\Gamma^{0}_{\underline{\hat{\chi}}2}>100\Gamma^{0}_{2},$
for sufficiently large $\mathcal{I}^{0}$ such that $\mathcal{I}^{0}a^{-\frac{1}{20}}<1$ and moreover choose the shift $b=0$ and the lapse $\Omega=1$ on the initial null hypersurface $H_{u_{\infty}}$. Then there exists a unique smooth solution $g$ of the vacuum Einstein equations in the region
$\mathcal{D}:=\big\{(u,\underline{u})\mid u_{\infty}\leq u\leq-a,\;\;0\leq\underline{u}\leq\epsilon\big\},$
satisfying the constraint and transport equations, with all geometric quantities $\psi\in\{\mbox{tr}\chi,\hat{\chi},\underline{\hat{\chi}},\mbox{tr}\underline{\chi},\eta,\underline{\eta},\alpha,\underline{\beta},\rho,\sigma\}$ remaining uniformly controlled by constants depending only on the initial data.
(b) Dynamical Existence of an apparent Horizon. Suppose, in addition to the bound 37-42, that the initial data along $\underline{H}_{0}$ satisfies the bound
(44) $\frac{3\pi(\text{Rad}(\widehat{\mathcal{M}}_{int})-\epsilon^{\frac{1}{2}})}{[\text{Rad}(\widehat{\mathcal{M}}_{int})]^{2}}+\frac{C(\Gamma^{0}_{2})}{a^{2}}<-\frac{2|u_{\infty}|\mbox{tr}\underline{\chi}(u_{\infty},0)}{a}+\frac{2}{a}\int_{u_{\infty}}^{a}|u^{{}^{\prime}}||\underline{\hat{\chi}}|^{2}(u^{{}^{\prime}},0)du^{{}^{\prime}}<\frac{3\pi}{\text{Rad}(\widehat{\mathcal{M}}_{int})}+\frac{C(\Gamma^{0}_{2})}{a^{2}},$
for a constant $C$ depending on the scale-invariant initial data norm $\Gamma^{0}_{2}$ controlling the Ricci coefficients $(\hat{\chi},\mbox{tr}\chi,\omega,\underline{\omega},\eta,\underline{\eta})$, uniformly for every direction along $\underline{u}=0$ and the radius $\text{Rad}(\widehat{\mathcal{M}}_{int})$ of the interior Cauchy slice at canonical double null time t=at=-a in the construction is prescribed to construct its Cauchy data, and the null expansion $\mbox{tr}\chi>0$ on $u=u_{\infty}$ compatible with the scaling 37-42, then the interior Cauchy slice $J^{+}(\widehat{\mathcal{M}}_{int})\cap\mathcal{M}_{t=-a+\epsilon}\subset\mathcal{M}_{t=-a+\epsilon}$ at canonical double null time $t=-a+\epsilon$ will contain an apparent horizon in its interior.
3.1. Main Ideas and Novelty
The principal analytical difficulty of the present work lies in the construction of the initial data, whether prescribed on a characteristic or a Cauchy hypersurface. The heuristics discussed in the preceding section suggest that the nonlinearities arising in the transport equations for the null Ricci coefficients contribute, in a delicate sense, only subdominant effects to the leading dynamics. Arbitrary initial data, however, fail to exhibit such cancellation structures, and in fact, by the celebrated small-data result of Christodoulou–Klainerman [17], one already knows that for sufficiently small perturbations of Minkowski data, no apparent horizon or trapped surface can form.
In the large-data regime, the situation is profoundly different: the system is fully nonlinear, no global small parameter exists, and long-time uniform control cannot, in general, be expected in the hyperbolic setting. As a result, one faces the possibility of geometric pathologies, including singularity formation and breakdown of the foliation. A central challenge, therefore, consists in constructing a hierarchy of large but controlled initial data $(g,k)$ that is compatible with a semi-global existence theorem for the Einstein vacuum equations, and whose evolution can be followed up to the onset of an apparent horizon.
From a structural viewpoint, the formation of a trapped surface is governed by the Raychaudhuri equation for the null expansion $\mbox{tr}\chi$ along the outgoing null direction:
@@ 3.1. Main Ideas and Novelty
In the traditional approach (cf. [16, 22, 1, 3]), one attempts to produce a negative expansion $\mbox{tr}\chi<0$ within a finite affine parameter time along the null generators. This mechanism is driven by the largeness of the incoming gravitational shear $|\widehat{\chi}|^{2}$, which acts as a source in (45).
@@
as introduced in Section 1.2. Our objective is to obtain a strictly positive value of $c$ on an appropriate boundary hypersurface, thus triggering the dynamical emergence of an apparent horizon from a previously horizon–free configuration. In the double–null gauge, one has the decomposition

### v4

@@ 2.4. Radius Comparison
(a) Semi–Global Existence of a Characteristic initial value problem: For every sufficiently large fixed constant $\mathcal{I}>0$, there exists $a_{0}=a_{0}(\mathcal{I})\gg 1$ such that the following holds. Let $a>a_{0}$ and let $N\in\mathbb{N}$ be sufficiently large. Suppose the initial data set is prescribed along the characteristic hypersurfaces $\underline{H}_{0}$, where $u_{\infty}<0$ is large in magnitude with
@@
for sufficiently large $\mathcal{I}^{0}$ and moreover dispersive data is prescribed on the initial null hypersurface $H_{u_{\infty}}$ compatible with the following smooth data
[…]
$\mbox{tr}\chi-\frac{2}{|u|_{\infty}}=O(a^{-\frac{1}{2}}|u_{\infty}|^{-1}),~\hat{\chi}=O(a^{-\frac{1}{2}}|u_{\infty}|^{-1}),~\mbox{tr}\underline{\chi}+\frac{2}{|u_{\infty}|}=O(|u_{\infty}|^{-2}),~\underline{\hat{\chi}}=O(a^{\frac{1}{2}}|u_{\infty}|^{-2}),$
$\omega=O(a^{\frac{1}{2}}|u_{\infty}|^{-2}),~\underline{\omega}=O(a^{\frac{1}{2}}|u_{\infty}|^{-3}),~(\eta,\underline{\eta})=O(a^{\frac{1}{2}}|u_{\infty}|^{-2})$
$\alpha=O(a^{-\frac{1}{2}}|u_{\infty}|^{-1}),~\beta=O(|u_{\infty}|^{-2}),~\underline{\beta}=O(a|u_{\infty}|^{-4}),~(\rho,\sigma)=O(a^{\frac{1}{2}}|u_{\infty}|^{-3}),~\underline{\alpha}=O(a^{\frac{3}{2}}|u_{\infty}|^{-5}),$
where $\gamma_{0}$ is the standard unit round metric in a topological 22-sphere. Then there exists a unique smooth solution $g$ of the vacuum Einstein equations in the region
$D_{a,1}:=\big\{(u,\underline{u})\mid u_{\infty}\leq u\leq-a,\;\;0\leq\underline{u}\leq 1\big\},$
satisfying the constraint and transport equations, with all geometric quantities $\psi\in\{\mbox{tr}\chi,\hat{\chi},\underline{\hat{\chi}},\mbox{tr}\underline{\chi},\eta,\underline{\eta},\alpha,\underline{\alpha},\underline{\beta},\rho,\sigma\}$ remaining uniformly controlled by constants depending only on the initial data.
(b) Local Cauchy Evolution: Fix integers s,Ns,~N large enough and let $a\gg 1$. Let $\mathcal{M}_{-a}$ be a smooth spacelike Cauchy hypersurface decomposed as
$\mathcal{M}_{-a}=\mathcal{M}_{1}\;\cup\;\mathcal{M}_{2}\cup\mathcal{M}_{ext},\qquad\partial\mathcal{M}_{1}=S_{-a,0},$
where $\mathcal{M}_{2}\subset D_{a,1}$ is generated by a canonical characteristic development from past null infinity and is foliated by a double–null optical function pair $(u,\underline{u})$ with $t=u+\underline{u}=-a$.
Let $(g,k)$ be smooth vacuum Cauchy data on a slightly bigger domain $\mathcal{M}^{1/a}_{int}:=\mathcal{M}_{1}\cup(\mathcal{M}_{2}\setminus(D^{{}^{\prime}}\cap\mathcal{M}_{2})$ with $D^{{}^{\prime}}\subset D_{a,1}$ and $D^{{}^{\prime}}$ is bounded by $u=-a-\frac{1}{a},\underline{u}=0,\underline{u}=1,$ and $u=u_{\infty}$, and assume that on $\widetilde{\mathcal{M}}_{1}\subset\mathcal{M}^{1/a}_{\mathrm{int}}$ the initial data satisfy the smallness bounds
(56) $\|\partial^{k+1}g(-a)\|_{L^{\infty}(\widetilde{\mathcal{M}}_{1})}\ \lesssim\ a^{-k-1},~\|\partial^{k}k(-a)\|_{L^{\infty}(\widetilde{\mathcal{M}}_{1})}\ \lesssim\ a^{-k-3/2},$
[…]
(57) $\|\partial^{k+1}g(-a)\|_{H^{s-1}_{ul}(\widetilde{\mathcal{M}}_{1})}\ \lesssim\ a^{-k-1},~\|\partial^{k}k(-a)\|_{H^{s-1}_{ul}(\widetilde{\mathcal{M}}_{1})}\ \lesssim\ a^{-k-3/2},~k\geq 0,$
where on the collar $\mathcal{M}^{1/a}_{int}\setminus\widetilde{\mathcal{M}}_{1}$ of thickness $O(\widetilde{\epsilon}a),~\widetilde{\epsilon}\in[\frac{1}{1000},\frac{1}{100}]$ the estimate reads
(58) $\|\partial^{k}k\|_{L^{\infty}(\mathcal{M}^{1/a}_{int}\setminus\widetilde{\mathcal{M}}_{1})}\leq Ca^{-\frac{3}{2}-k},$
[…]
(60) $\|\partial^{k}k\|_{H^{s-1}_{\mathrm{ul}}(\mathcal{M}^{1/a}_{int}\setminus\widetilde{\mathcal{M}}_{1})}\leq Ca^{-k-3/2},~k\geq 0,$
and $(k,\partial g)$ smoothly interpolates between the data on $\widetilde{\mathcal{M}}_{1}$ and the data on $\Sigma_{int}=\mathcal{M}_{-1}\cap D_{a,1}$ in spacetime harmonic coordinates. Assume moreover that the harmonic gauge constraints and the Einstein constraint equations hold initially.
Assume that $(g,k)$ is smoothly matched across $S_{-a,0}$ with the Cauchy data induced on $\mathcal{M}_{2}$ by the characteristic development $D_{a,1}$, in the sense that:
(1) the induced metric and second fundamental form on $S_{-a,0}$ agree when computed from $(\mathcal{M}_{1},g)$ and from the characteristic spacetime $(\mathcal{M}_{\mathrm{ext}}\cap D_{a,1},g_{\mathrm{char}})$,
(2) all tangential derivatives up to order $N$ coincide on $S_{-a,0}$.
[…]
$m_{\mathrm{ADM}}\sim a^{\frac{1}{2}},$
and angular momentum $J=O(a)$, preserving the vacuum constraints and the above Sobolev bounds. Let
[…]
$\varepsilon=\varepsilon\!\left(N,\,\mathrm{Rad}(\mathcal{M}_{1}),\,\|\tilde{g}\|_{H^{N+1}_{ul}},\,\|\tilde{k}\|_{H^{N}_{ul}}\right)>0$
such that the vacuum Einstein equations admit a unique local Cauchy development
$(\mathcal{M}\times[-a,-a+\varepsilon],\mathsfbf g)$
[…]
  1. (1)
    $\mathsfbf g\in C^{0}_{t}H^{N+1}_{x}\cap C^{1}_{t}H^{N}_{x}$;
  2. (2)
    the harmonic gauge constraints propagate on $a,-a+\varepsilon]$;
  3. (3)
    the solution coincides with the characteristic development $D_{a,1}$ in the overlap region.
Moreover, for all sufficiently large $a$, the lifespan $\varepsilon$ can be chosen uniformly, and in particular one may take
$\varepsilon=\frac{3}{4}.$
(c) Dynamical Formation of a MOTS: Suppose, in addition to the assumption of part $(a)$ and $(b)$, the initial Riemannian metric $g$ on $\mathcal{M}_{1}$ has $H$-radius $\text{Rad}(\mathcal{M}_{1})\gg 1$ and set $\text{Rad}(\mathcal{M}_{1})=\frac{3\pi}{4}(a-1)+O(a^{-1}),~a\gg 1$ and the initial data on $\underline{H}_{0}$ also verify the following isotropic condition
(61) $\frac{17}{9a}<\int_{u_{\infty}}^{-a}|u^{\prime}||\underline{\hat{\chi}}|^{2}(u^{\prime},\epsilon)\,du^{\prime}+\int_{u_{\infty}}^{-a}\frac{1}{|u^{\prime}|^{2}}\int_{u_{\infty}}^{u^{\prime}}|u^{\prime\prime}|^{2}|\underline{\hat{\chi}}|^{2}(u^{\prime\prime},\epsilon)\,du^{\prime\prime}\,du^{\prime}<\frac{19}{9a},$
then the Cauchy slice $\mathcal{M}_{-a}$ will not contain a MOTS and the interior Cauchy slice $J^{+}(\mathcal{M}^{1/a}_{int})\cap\mathcal{M}_{t=-a-1/a+\epsilon}\subset\mathcal{M}_{t=-a-1/a+\epsilon}$ at canonical double null time $t=-a-1/a+\epsilon,~\epsilon=\frac{3}{4}$ will contain a MOTS in its interior.
Remark 3.
Notice that given $\underline{\hat{\chi}}$ on $\underline{H}_{0}$, its geometry is determined once boundary condition in $S_{u_{\infty},0}$ is given. More precisely along with the conjugate shear, (a) the conformal class/metric on the corner sphere $S_{u_{\infty},0}$ and (b)a gauge normalization, for example: $\Omega$ fixed on $S_{u_{\infty},0}$, together with a choice of the shift vector $b$ along the initial null hypersurface $\underline{H}_{0}$. In addition, the decay rate for the Weyl curvature components are fixed by the peeling property [26], while the connection coefficients are less fundamental and dependent on the choice of frame.
Remark 4.
The scaling $a^{-3/2}$ is critical for compatibility with the Bel–Robinson energy hierarchy on time slabs of size $O(1)$.
The uniformity of $\varepsilon$ relies on the effective point-wise smallness of the initial k,|N1|k,|N-1|, and $\nabla N$; largeness is absorbed in the largeness of the Hawking mass of the spheres foliating the characteristic development.
Remark 5.
Notice that the condition on the strict upper bound on the Schoen-Yau radius of the interior $\mathcal{M}_{1}$ is very flexible in terms of the choice of the corresponding metrics in light of constructions of [44]. On the other hand, $a^{\frac{1}{2}}$ is roughly related to the ADM parameters of the initial slice $\mathcal{M}_{-a}$. This is precisely the idea of a large mass contained in an isotropically (in terms of the $H$-radius or Schoen-Yau radius) large domain. In particular, the interior domain $\mathcal{M}_{1}$ needs to be uniformly thick or isotropically large measured in terms of $H$-radius.
Remark 6.
Notice the vital importance of the interior estimate $|\partial g|=O(a^{-1})$ (in harmonic coordinates) in the interior $\widetilde{\mathcal{M}}_{1}$. This implies through fundamental theorem of calculus $|g-\delta|=O(1)$ in the interior which leads to the length estimate $|d_{g}(l)-d_{\delta}(l)|=O(a)$ on $\widetilde{\mathcal{M}}_{1}$ of a curve $l$ of $\delta$-length $O(a)$. This is necessary for $H$-radius to be $\text{Rad}(\mathcal{M}_{1})=\frac{3\pi}{4}(a-1)+O(a^{-1})$ in the interior-in particular the largeness of the interior $H$-radius originates form the largeness of the weighted entity $|a\partial g|$. Notice that if instead one had $|\partial g|=O(a^{-1-\delta})$ in the interior, then similarly one would have obtained $|g-\delta|=O(a^{-\epsilon})$ and $|d_{g}(l)-d_{\delta}(l)|=O(a^{1-\epsilon})$ which would be insufficient to satisfy the condition $\text{Rad}(\mathcal{M}_{1})=\frac{3\pi}{4}(a-1)+O(a^{-1})$ due to $a^{-\epsilon}$ loss.
2.5. Comparison with Previous Studies and Novelty
Following the singularity theorem of Penrose [34], the classical mechanism guaranteeing the presence of a black hole region requires the prior existence of a closed trapped surface. For a considerable period, the only available route to such a surface was to impose its existence directly at the level of the initial data set. This requirement is itself highly nontrivial, since the trapped surface condition is a nonlinear, fully geometric constraint involving both the intrinsic and extrinsic geometry of the initial slice.
A decisive advance at the level of initial data was achieved by Schoen–Yau [43], who established the first general existence theorem for marginally outer trapped surfaces (MOTS) under suitable geometric and energy conditions. Subsequently, Yau [44] discovered a substantially stronger principle: the existence of a MOTS can be forced purely by boundary geometry. More precisely, certain quantitative boundary convexity conditions imply the existence of a MOTS in the interior, independent of any positivity assumption on the matter density (and in particular allowing negative energy densities). This boundary–driven mechanism revealed a fundamentally new geometric effect and indicated that trapped surface formation is not exclusively tied to bulk matter concentration. This is the primary motivating point of this current study (See [7, 8] for studies related to dynamic and trapping horizon, later works on Jang’s equation and its relationship with MOTS).
These results naturally lead to a deeper dynamical question: are MOTSs genuinely evolutionary objects, in the sense that they arise from the Einstein evolution of regular initial data containing no trapped or marginally trapped surfaces? From both mathematical and physical perspectives, this issue is central. The geometric definition of a black hole region in general relativity derives its significance from its predictive and observational content; it must correspond to objects that can form dynamically from physically admissible configurations. In this sense, dynamical trapped surface formation constitutes a stringent consistency requirement linking the analytic theory of the Einstein equations with the physical interpretation of black holes. This is particularly relevant for astrophysical scenarios such as supermassive black holes, whose formation is not adequately modeled by short–pulse collapse mechanisms and therefore demands a large–scale, genuinely dynamical geometric theory of trapped surface formation.
The first results along this direction were obtained by Christodoulou for the Einstein equations coupled to a massless scalar field in spherical symmetry. Through a series of works [18], [16], [17], and [20], Christodoulou managed to not only prove trapped surface formation, but to understand the picture of gravitational collapse in its entirety for the given model and under the given symmetry. The breakthrough in the absence of symmetry came in [19] by the same author. In this work, Christodoulou introduced a hierarchy of small and large components in the initial data which (almost) persists under the evolution of the Einstein equations. He termed his method the short pulse method. After Christodoulou, another breakthrough work by Klainerman-Rodnianski [25] reduces the size of Christodoulou’s work from about 600 to approximately 120 pages by using a slightly different hierarchy. Moreover, it reduces the number of derivatives of curvature required to prove semi-global existence from two to one through refined trace estimates. A substantial extension of the result of Christodoulou (which required a uniform condition along all null geodesic generators instead) was executed by Klainerman-Luk-Rodnianski [24], proving a fully genuine anisotropic criterion for the formation of trapped surfaces in vacuum. More precisely, they provide local conditions on null data, concentrated in a neighborhood of a short null geodesic segment (possibly flat in all other directions) whose future development contains a trapped surface.
A few years later, An [3] introduces the signature for decay rates $s_{2}$ on his way to proving an extension of [25] from a finite region to a region close to past null infinity-this method proved to be very efficient in handling large data problems in general. In 2014, An and Luk [6] proved the first scale-critical trapped surface formation criterion for the vacuum equations in the absence of symmetry. While Christodoulou’s data in [19] were large in $\dot{H}^{1}(\mathbb{R}^{3})$, An and Luk give data which only have to be large in $\dot{H}^{\frac{3}{2}}(\mathbb{R}^{3})$, which is a scale-critical norm for the initial data. Taking advantage of the scale criticality in [6], An [2] constructs initial data that give rise not merely to trapped surfaces, but an apparent horizon, a smooth 3-dimensional hypersurface consisting of marginally outer trapped surfaces. In 2019, An [1] produces a 55-page proof of trapped surface formation for the vacuum equations, making use of the signature for decay rates and obtaining an existence result from a region close to past null infinity. In [4], An and Athanasiou extended [1] to the case of the Einstein-Maxwell system. Several other studies exist in the context of Einstein-Yang-Mills [12], Einstein-Vlasov system [23, 9, 10, 13], Einstein-Scalar field system, and Einstein-Spinor field system. Recently, [15], in a fundamental study, provided a short proof of the formation of a trapped surface in geodesic foliation.
The principal analytical difficulty of the present work lies in the prescription of the initial data, whether prescribed on a characteristic or a Cauchy hypersurface. One of the main novelties of this study and the difference with the previous studies is the delicate nature (cancellation) of the estimates associated with the transport equation. Arbitrary initial data, however, fail to exhibit such cancellation structures, and in fact, by the celebrated small-data result of Christodoulou–Klainerman [19], one already knows that for sufficiently small perturbations of Minkowski data (in global sense), no MOTS or trapped surface can form.
In the large-data regime, the situation is profoundly different: the system is fully nonlinear, no global small parameter exists, and long-time uniform control cannot, in general, be expected in the hyperbolic setting. As a result, one faces the possibility of geometric pathologies, including singularity formation and breakdown of the foliation. A central challenge, therefore, consists in constructing a hierarchy of large but controlled initial data $(g,k)$ that is compatible with a semi-global existence theorem for the Einstein vacuum equations, and whose evolution can be followed up to the onset of an MOTS.
From a structural viewpoint, the formation of a trapped surface is governed by monotonicity property of the Raychaudhuri equation for the null expansion $\mbox{tr}\chi$ along the outgoing null direction:
[…]
@@ 2.5. Comparison with Previous Studies and Novelty
In the previous approach (cf. [18, 25, 1, 3]), one attempts to produce a negative expansion $\mbox{tr}\chi<0$ within a finite affine parameter time along the null generators. This mechanism is driven by the largeness of the incoming gravitational shear (radiation) $|\widehat{\chi}|^{2}$, which acts as a source in (62).
@@
as introduced in Section 2.1 and considered by Schoen-Yau [43] in the context of matter sourced gravity and by [44] in a more general framework that included the case of negative energy density (and pure vacuum in particular). Our objective is to obtain a strictly positive value of $c$ on an appropriate boundary hypersurface of an isotropically large (appropriately defined) domain (i.e., the $H$-radius or the Schoen-Yau radius is large), thus triggering the dynamical emergence of an MOTS from a previously horizon-free configuration. In the double-null gauge, one has the decomposition